### v1

@@ 2 Preliminaries
The following lemma is standard (see e.g. [24, Corollary 9]).
@@ 3 Grid graphs and extensions
We next argue that there is a subpath $P$ of $T$ which contains the nodes $x(1),\dots,x(b)$ in this particular order. There are two alternatives to exclude.
@@
  • Suppose that for some $1\leq i_{1}<i_{2}<i_{3}\leq b$, the node $x(i_{3})$ lies on the path between $x(i_{1})$ and $x(i_{2})$ in $T$. Within $D_{n,m}$, there are $n$ vertex disjoint paths from $X_{i_{1}}$ to $X_{i_{2}}$ that do not pass through $X_{i_{3}}$. In particular, $B_{x(i_{3})}$ now must contain one vertex for each such path, plus the $n$ vertices in $X_{i_{3}}$. But then $|B_{x(i_{3})}|\geq 2n$, contradicting the fact that all bags have size at most 2n12n-1.
    A similar argument applies to exclude the possibility that $x(i_{1})$ lies in between $x(i_{2})$ and $x(i_{3})$.
  • Suppose that for some $1\leq i_{1}<i_{2}<i_{3}\leq b$, there is a node $y\in V(T)$ which has three internally vertex-disjoint paths to $x(i_{1}),x(i_{2})$ and $x(i_{3})$. Then $|B_{y}|\geq 2n$ (contradicting the width again) since it must contain at least one vertex of each of the following $2n$ vertex-disjoint paths: the $n$ paths from $X_{i_{1}}$ to $X_{i_{2}}$ which keep their first coordinate fixed and the $n$ paths from
    $X_{i_{2}}+(0,1)=\{(x,y+1):(x,y)\in X_{i_{2}}\}$
    to $X_{i_{3}}$ which keep their first coordinate fixed.
So $T$ has a subpath $P$ containing the nodes $x(1),\dots,x(b)$ in this particular order. Next, we show that
$v_{0}\in B_{x(n\cdot n^{2}+1)}\cap B_{x((n+1)n^{2}+1)}\cap\dots\cap B_{x(b-n\cdot n^{2}+1)}.$ (1)
@@
Suppose first towards a contradiction that $v_{0}\not\in B_{x(a)}$ for all $a<j$ yet $v_{0}\in B_{x(j)}$ for some $j\in\{n\cdot n^{2}+1,\dots,b\}$. For $\ell\in\{0,1,\dots,n-1\}$, we choose a vertex $x_{\ell}\in X_{\ell n^{2}+1}$ and define a path $P_{\ell}$ within $J_{\ell n^{2}+1}\cup\{v_{0}\}$ running from $v_{0}$ to $x_{\ell}$ via the neighbour $(1,y_{\ell n^{2}+1})$ of $v_{0}$. The paths $P_{\ell}$ intersect only in $v_{0}$ and are disjoint from $X_{j}$. Recall that the vertices in $X_{a}$ are by definition in the bag $B_{x(a)}$. Since $v_{0}\not\in B_{x(a)}$ for all $a<j$ and $v_{0}\in B_{x(j)}$, for each $\ell\in\{0,1,\dots,n-1\}$, $B_{x(j)}$ must either contain $x_{\ell}$, or $B_{x(j-1)}\cap B_{x(j)}$ is a separator between $x_{\ell}$ and $v_{0}$. Either way, $B_{x(j)}$ must contain a vertex of $P_{\ell}\setminus\{v_{0}\}$. Since the $n$ vertices from $X_{j}$ are also present in $B_{x(j)}$, using the disjointness constraints, we obtain $2n$ different vertices in $B_{x(j)}$, contradicting the upper bound on the width. A similar argument can be used with the neighbours $(1,y_{b-n^{2}+1}),\dots,(1,y_{b-n\cdot n^{2}+1})$ to exclude the possibility that $v_{0}\not\in B_{x(a)}$ for all $a>j$ yet $v_{0}\in B_{x(j)}$, for all $j\in\{1,\dots,b-n\cdot n^{2}+1\}$.
We showed above that if $v_{0}\in B_{x(j)}$ for some $j\in[b]$, then it is also in $B_{x(a)}$ for all $a\in\{j,\dots,b-n\cdot n^{2}+1\}\cup\{n\cdot n^{2}+1,\dots,j\}$. The remaining case in which (1) fails is when $v_{0}\not\in B_{x(j)}$ for all $j\in[b]$. Then there is a unique $j$ such that $x(j)$ is closest among $x(1),\dots,x(b)$ to a bag containing $v_{0}$. The same argument as above can now be applied, since $B_{x(j-1)}\cap B_{x(j)}$ (or $B_{x(j)\cap B_{x(j+1)}}$) will still form a separator in this case between $v_{0}$ and $x_{\ell}$.
@@
In conclusion, we established that $v_{0}$ is present in at least $b-2(n+1)n^{2}=n^{4}-2(n+1)n^{2}$ bags while its degree is $b/n^{2}=n^{2}$. For $n$ sufficiently large, the spread of $v_{0}$ is at least $n^{4}-2(n+1)n^{2}>\frac{1}{2}n^{2}(n^{2}+1)=\frac{1}{2}n^{2}d(v)$.
@@ 4 Well-behaved spread using thrice the width
Suppose that $|V(G)|>(3+8/b)k$ and that the lemma has been shown already for graphs on fewer vertices. Let $S$ be given as in the lemma and give all vertices in $S$ weight $1$ and the remaining vertices weight 0. We apply Lemma 2.1 with this weight function to find a $t$ for which for each component $C$ of $G-B_{t}$, we find $|C\cap S|\leq|S|/2$.
Let $X=B_{t}$. For each component $C$ of GXG-X, we add vertices to a set $S_{C}$ and define a marking $m_{C}^{\prime}:S_{C}\to[b]$ as follows:
  • For all vertices $v$ in $S$ marked 1 by $m^{\prime}$ (which may be in $X$) that have a neighbour in $C\setminus S$, we add $v$ and a neighbour of $v$ in $C\setminus S$ to $S_{C}$, give $v$ mark $b$ and also give the neighbour mark $b$.
  • All other vertices $s$ in $S\cap C$ that have a neighbour in $C\setminus S$, we add with mark $m^{\prime}(s)-1$. (Here $m^{\prime}$ was given in the lemma statement.)
  • All vertices $x$ in $X\setminus S$ with a neighbour in $C\setminus S$, we add with mark b1b-1.
@@
In order to ensure the marking $m^{\prime}_{C}$ for $S_{C}$ satisfies the assumptions in the induction hypothesis/lemma statement, we can decrease some of the marks if needed. By definition, at this point at most $(1/b)(2+8/b)k$ vertices in $S_{C}$ are marked $i$ for some $i\in[b-2]$, since only vertices that were previously marked $i+1$ in $S$ by $m^{\prime}$ can receive mark $i$ for $i\in[b-1]$ by $m_{C}^{\prime}$. Moreover, at most
@@
vertices can ever receive a mark $\leq b-1$. This means that if too many vertices received mark b1b-1 at this point, we can safely decrease some of their marks (‘increasing their priority’).
@@
Let $c=1+2t$. We will prove that every graph $G$ has a tree-decomposition $(T,(B_{t})_{t\in V(G)})$ of width at most $c\operatorname{tw}(G)$ and average spread
@@
In [24, Lemma 23] it is shown that such a quotient yields a tree decomposition $(F,(B_{x}^{\prime})_{x\in V(F)})$ of $G$, for any tree $F$ with vertex-set $\{1,\dots,m\}$, rooted at vertex $1$, where for $i\in\{2,...,m\}$, the parent of $i$ is any number $j\in\{1,\dots,i-1\}$ such that $r_{i}\in V(T_{j})$.
@@ 6 Conclusion
We remark that for the $(n\times n)$-grid, it is already highly non-trivial to prove the rankwidth is n1n-1 (as shown by Jelínek [18]) or that the cliquewidth is $n+1$ (as shown by Golumbic and Rotics [16]).

### v2

@@ 1 Introduction
The parameter treespan can be defined in terms of the occupation time in a graph searching game [15] and is polynomially tied to domino treewidth (see [19, Theorem 54]).
@@ 2 Preliminaries
The following lemma is standard (see e.g. [27, Corollary 9]).
@@ 3 Grid graphs and extensions
Let $P$ denote the path in $T$ between $x(1)$ and $x(b)$. For each $a\in[b]$, we let $y(a)$ denote the node in $P$ closest to $x(a)$. (For example, $y(1)=x(1)$ and $y(b)=x(b)$.)
Claim 1.
The nodes $y(a)$ are distinct and appear in the order $y(1),\dots,y(b)$ in $P$.
Proof 3.7.
@@
First, we exclude the option of the existence of $1\leq i_{1}<i_{2}<i_{3}\leq b$ such that the node $x(i_{3})$ lies on the path between $x(i_{1})$ and $x(i_{2})$ in $T$. Within $D_{n,m}$, there are $n$ vertex disjoint paths from $X_{i_{1}}$ to $X_{i_{2}}$ that do not pass through $X_{i_{3}}$. In particular, $B_{x(i_{3})}$ now must contain one vertex for each such path, plus the $n$ vertices in $X_{i_{3}}$. But then $|B_{x(i_{3})}|\geq 2n$, contradicting the fact that all bags have size at most 2n12n-1. A similar argument applies to exclude the possibility that $x(i_{1})$ lies in between $x(i_{2})$ and $x(i_{3})$.
Since the $x(a)$ must be distinct (by the width assumption), the claim above shows that for each $a\in\{2,\dots,b-1\}$, the node $y(a)$ must lie strictly between $y(1)$ and $y(b)$ on $P$.
We next show that all $y(a)$ are distinct. Suppose that there are $1<a<a^{\prime}<b$ with $y=y(a)=y(a^{\prime})$. Then $B_{y}$ is a vertex separator between $2n$ vertex-disjoint paths, $n$ from $X_{1}$ to $X_{a}$ and $n$ from $X_{a^{\prime}}$ to $X_{b}$. This contradicts the width assumption since that would imply that $|B_{y}|\geq 2n$.
Finally, we exclude the option that for some $1<a<a^{\prime}<b$, the nodes appear in the wrong order: $y(1),y(a^{\prime}),y(a),y(b)$. Again, both $y(a)$ and $y(a^{\prime})$ are now a vertex separator between $2n$ vertex-disjoint paths, $n$ from $X_{1}$ to $X_{a}$ and $n$ from $X_{a^{\prime}}$ to $X_{b}$.
So the nodes $y(1),\dots,y(b)$ are distinct and appear in this particular order.
Next, we show that
$v_{0}\in B_{y(n\cdot n^{2}+1)}\cap B_{y((n+1)n^{2}+1)}\cap\dots\cap B_{y(b-n\cdot n^{2}+1)}.$ (1)
Let $z(1),\dots,z(b^{\prime})$ denote all nodes of $P$ chosen in order with $z(1)=y(1)$. If the equation above fails, then there are two options. First, if $v_{0}\not\in B_{z}$ for all $z\in V(P)$, then let $z(i_{0})$ denote the (unique) closest node in $P$ to a bag containing $v_{0}$. Second, if $v_{0}\in B_{z}$ for some $z\in V(P)$, then let $z_{\min}$ (respectively $z_{\max}$) be the $z\in[b^{\prime}]$ closest to $y(1)$ (respectively $y(b)$) such that $v_{0}\in B_{z}$. If $z_{\min}$ appears before $y(n\cdot n^{2}+1)$ (or is equal to it) and $z_{\max}$ appears after $y(b-n\cdot n^{2}+1)$ (or is equal to it), then (1) holds. If $z_{\min}$ appears strictly after $y(n\cdot n^{2}+1)$ on $P$ then we define $z(i_{0})=z_{\min}$. Otherwise, $z_{\max}$ must appear strictly before $y(b-n\cdot n^{2}+1)$ and we define $z(i_{0})=z_{\max}$.
Either way, $z(i_{0})$ is defined such that it either appears strictly after $y(n\cdot n^{2}+1)$ on $P$ and is the ‘first’ node closest to a bag containing $v_{0}$, or strictly before $y(b-n\cdot n^{2}+1)$ and is the ‘last’ such node. Our argument will be analogous in both cases and we will assume henceforth that the former holds.
@@
Let $j\in[a]$ be the smallest such that $y(j)=z(i_{0})$ or $y(j)$ appears after $z(i_{0})$. By assumption, $j>n\cdot n^{2}+1$. For $\ell\in\{0,1,\dots,n-1\}$, we choose a vertex $x_{\ell}\in X_{\ell n^{2}+1}$ and define a path $P_{\ell}$ within $J_{\ell n^{2}+1}\cup\{v_{0}\}$ running from $v_{0}$ to $x_{\ell}$ via the neighbour $(1,y_{\ell n^{2}+1})$ of $v_{0}$. The paths $P_{\ell}$ intersect only in $v_{0}$ and are disjoint from $J_{j-1}\cup J_{j}\cup J_{j+1}$. There are $n$ vertex-disjoint paths from $X_{j-1}$ to $X_{j+1}$ entirely contained in $J_{j-1}\cup J_{j}\cup J_{j+1}$ (that keep their coordinate fixed) and since $z(i_{0})$ lies on the path between x(j1)x(j-1) and $x(j+1)$ (by definition of $j$), $B_{z(i_{0})}$ contains at least one vertex from each of these $n$ paths. Moreover, by definition of $i_{0}$, $v_{0}\not\in B_{z(i_{0}-1)}$ and $v_{0}\in B_{z(i_{0})}$. This shows that $B(z(i_{0})\cap B_{z(i_{0}-1)}$ is a vertex separator between $X_{\ell}$ and $\{v_{0}\}$ which does not contain $v_{0}$, and so for each $\ell\in\{0,1,\dots,n-1\}$, $B_{z(i_{0}-1)}\cap B_{z(i_{0})}$ must contain a vertex of $P_{\ell}\setminus\{v_{0}\}$. Using the disjointness constraints, we obtain $2n$ different vertices in $B_{z(i_{0})}$, contradicting the upper bound on the width.
@@
In conclusion, we established that $v_{0}$ is present in at least $b-2(n+1)n^{2}=n^{4}-2(n+1)n^{2}$ bags while its degree is $b/n^{2}=n^{2}$. For $n$ sufficiently large, the spread of $v_{0}$ is at least $n^{4}-2(n+1)n^{2}>\frac{1}{2}n^{2}(n^{2}+1)=\frac{1}{2}n^{2}d(v_{0})$.
@@ 4 Well-behaved spread using thrice the width
Suppose that $|V(G)|>(3+8/b)k$ and that the lemma has been shown already for graphs on fewer vertices. Let $S,m^{\prime}$ be given as in the lemma statement. We define a weight function $w$ on $G$ by giving all vertices in $S$ weight $1$ and the remaining vertices weight 0. We apply Lemma 2.1 on $(G,w)$ using any tree decomposition for $G$ of optimal width, in order to find a vertex set $X$ with $|X|\leq k=\operatorname{tw}(G)+1$ such that each component $C$ of GXG-X satisfies $|C\cap S|\leq|S|/2$.
For each component $C$ of GXG-X, we add vertices to a set $S_{C}$ and define a marking $m_{C}^{\prime}:S_{C}\to[b]$ as follows:
  • For all vertices $v$ in $S$ marked 1 by $m^{\prime}$ (which may be in $X$) that have a neighbour in $C\setminus S$, we add $v$ and a fixed neighbour $u$ of $v$ in $C\setminus S$ to $S_{C}$ and set $m^{\prime}_{C}(v)=m^{\prime}_{C}(u)=b$.
  • All other vertices $s\in S\cap C$ with a neighbour in $C\setminus S$ are added to $S_{C}$ with mark $m^{\prime}_{C}(s)=m^{\prime}(s)-1$.
  • All vertices $x\in X\setminus S$ with a neighbour in $C\setminus S$ are added to $S_{C}$ add with mark $m_{C}(x)=b-1$.
@@
In order to ensure the marking $m^{\prime}_{C}$ for $S_{C}$ satisfies the assumptions in the induction hypothesis/lemma statement, we can decrease some of the marks if needed. By definition, at this point at most $(1/b)(2+8/b)k$ vertices in $S_{C}$ are marked $i$ for some $i\in[b-2]$, since only vertices that were previously marked $i+1$ in $S$ by $m^{\prime}$ can receive mark $i$ for $i\in[b-2]$ by $m_{C}^{\prime}$. Moreover, at most
@@
vertices can ever receive a mark $\leq b-1$. (The last step from this displayed equation follows from the previous displayed equation.) This means that if too many vertices received mark b1b-1 at this point, we can safely decrease some of their marks (‘increasing their priority’).
@@
Let $c=1+2t$. We will prove that every graph $G$ has a tree-decomposition $(T,(B_{t})_{t\in V(T)})$ of width at most $c\operatorname{tw}(G)$ and average spread
@@
In [27, Lemma 23] it is shown that such a quotient yields a tree decomposition $(F,(B_{x}^{\prime})_{x\in V(F)})$ of $G$, for any tree $F$ with vertex-set $\{1,\dots,m\}$, rooted at vertex $1$, where for $i\in\{2,...,m\}$, the parent of $i$ is any number $j\in\{1,\dots,i-1\}$ such that $r_{i}\in V(T_{j})$.
@@ 6 Conclusion
We remark that for the $(n\times n)$-grid, it is already highly non-trivial to prove the rankwidth is n1n-1 (as shown by Jelínek [20]), that the cliquewidth is $n+1$ (as shown by Golumbic and Rotics [17]) or that its edge-bandwidth is 2n12n-1 (as shown by Pikhurko and Wojciechowski [22]).